OpenStudy (anonymous):

What is the solution of the system defined by y = -x + 5 and 5x + 2y = 14? x=1/3, y=11/7 x=4/3, y=11/3 x=5/11, y=0 x=4/3, y=1 all i want is the answer,i need to get this done asap

OpenStudy (anonymous):

Can u write the second equation in terms of 'y'

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

Make 'y' the subject

OpenStudy (anonymous):

I dont know what that means :/

OpenStudy (anonymous):

like this 'y= (14-5x)/2'

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

I'll help

OpenStudy (anonymous):

ok thanks,but i have 3 min left

OpenStudy (anonymous):

U in exam

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Naughty!!!

OpenStudy (anonymous):

i have to get 4/5 to pass and im at 4/5 so i cant get this wrong

OpenStudy (anonymous):

this one question says if i stay in my school or not

OpenStudy (anonymous):

Last option!

OpenStudy (anonymous):

*is correct

OpenStudy (anonymous):

it was wrong T-T

OpenStudy (anonymous):

Ahh sorry second one!

OpenStudy (anonymous):

i have one makeup answer :D

OpenStudy (anonymous):

*makeup

OpenStudy (anonymous):

simplify x = 2.8 + 0.2y

OpenStudy (anonymous):

soo x=1.4+0.1y?

OpenStudy (anonymous):

Leave it!

OpenStudy (anonymous):

U in which grade

OpenStudy (anonymous):

im in 9 going into 10

OpenStudy (anonymous):

9! great i spent two in year 9!! U should do it tooo

OpenStudy (anonymous):

i go to a state tech school,if i fail i get sent back to my sending school

OpenStudy (anonymous):

Skipping one question! Nah man don't worry

OpenStudy (anonymous):

i cant though,i had to get 4/5 of the EOS test to determine if i stay,i already got one wrong so im at 4/5...3/5 is kicked out...but i got the make-up question right :D

OpenStudy (anonymous):

3 min over!

OpenStudy (anonymous):

i know,when i got it wrong i got an extra 10,i didnt know i would get the extra time tho

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

thanks for all your help.

OpenStudy (anonymous):

Np!